A crucible containing an unknown mass of magnesium weighs 37,63 if the mass of the empty crucible was found to be 32,83.determine the number of moles of magnesium

1
Expert's answer
2022-03-28T17:18:04-0400

m(Mg) = 37.63 - 32.83 = 4.8 g

n(Mg) = m(Mg)/Mr(Mg) = 4.8 g / 24 g/mol = 0.2 mol